OpenStudy (anonymous):

HELP MEEEE! What is the center of the circle that has a diameter whose endpoints are (2, 7) and (-6, -1)? (4, 3) (-2, 4) (-2, 3)

OpenStudy (anonymous):

@jacob1233 ? want to help?

OpenStudy (anonymous):

@Jadeishere this one pretty please?

OpenStudy (anonymous):

sure lets see here...

OpenStudy (anonymous):

and this one... Find the distance between the points (-4, -5) and (3, -1).

OpenStudy (anonymous):

choices are \[\sqrt{65} or \sqrt{37} or \sqrt{17}\]

OpenStudy (jadeishere):

Hold on i'm going to give you a formula to use real quick >.<

OpenStudy (anonymous):

kk

OpenStudy (jadeishere):

\[\frac{ x _{2} + x _{1}}{ 2 } and \frac{ y _{2} + y _{1} }{ 2 } = midpoints\]

OpenStudy (anonymous):

sorry I'm swamping you guys... just a little desperate right now! Factor the following polynomial completely. -x 2 y 2 + x 4 + 4y 2 - 4x 2 choices: (x - 2)(x - 2)(x + y)(x - y) (x + 2)(x + 2)(x + y)(x - y) (x + 2)(x - 2)(x + y)(x - y)

OpenStudy (jadeishere):

Okay hold on, seriously. You're going to confuse yourself on which questions we're helping you with if you go too fast. One at a time, please

OpenStudy (anonymous):

oops. sorry.

OpenStudy (anonymous):

your first answer for the first question is (4,3) i think

OpenStudy (jadeishere):

Yes, it is.. Did you use that formula?

OpenStudy (anonymous):

me? no i just had to think for a sec but the formula would help.. but i didn't use it

OpenStudy (jadeishere):

Lol okay

OpenStudy (anonymous):

the formula was for which one? first question or second?

OpenStudy (anonymous):

the last question answer is 1st factor (x²): = - x²y² + x⁴ + 4y² - 4x² = x⁴ - 4x² + - x²y² + 4y² = x²(x² - 4) - y²(x² - 4) = (x² - y²)(x² - 4) = (x + y)(x - y)(x + 2)(x - 2) Answer: c. (x + 2)(x - 2)(x + y)(x - y)

OpenStudy (jadeishere):

First, the formula wouldn't help with the second. Th second you're supposed to use a pathagorean theorem so a^2 + b^2 = c^2. i need a graphing tool...

OpenStudy (anonymous):

lol...

OpenStudy (jadeishere):

hold on i'll use a paper thingy

OpenStudy (anonymous):

ok. you guys are winning serious brownie points right now.

OpenStudy (anonymous):

if you are ever in Grand Cayman, Cayman Islands hit me up

OpenStudy (anonymous):

hahaa, and i just dont really know what you are asking for the second question? and i will if im ever there

OpenStudy (jadeishere):

so the distance between (-4, -5) and (3, -1) is sqrt(65)

OpenStudy (jadeishere):

>.< Sorry i had to graph it

OpenStudy (anonymous):

yea. that ^ no problem its fine

OpenStudy (jadeishere):

One of my goals is to travel around the world, not the big countries like France or Brazil, but all of the small ones that go unseen >.<

OpenStudy (anonymous):

then come to Grand Cayman.

OpenStudy (jadeishere):

That's what I meant, so there's a chance i'll go there >.<

OpenStudy (anonymous):

want to help me with another one? only a few more before we get my score

OpenStudy (anonymous):

cool!

OpenStudy (jadeishere):

Okay >.<

OpenStudy (anonymous):

choices are: -5n^9 -5n^4 5n^4

OpenStudy (anonymous):

okay

OpenStudy (jadeishere):

brb

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Volleygirl is there a picture?

OpenStudy (anonymous):

of the problem? yes in the attached file

OpenStudy (anonymous):

okay sorry i didnt see it

OpenStudy (anonymous):

no problem.

OpenStudy (jadeishere):

-5n^4

OpenStudy (anonymous):

yeah i think ur right

OpenStudy (jadeishere):

Lol >.<

OpenStudy (anonymous):

you 2 are amazing.

OpenStudy (jadeishere):

Ha, funny >.< I'm merely good

OpenStudy (anonymous):

naaa ur to kind we are just here to help

OpenStudy (anonymous):

ur amazing so hush

OpenStudy (jadeishere):

(60 sqrt3)

OpenStudy (anonymous):

choices for this one are: \[60\sqrt{5}...60\sqrt{3}... or... 50\sqrt{3}\]

OpenStudy (anonymous):

wow. ur fast

OpenStudy (jadeishere):

Lol xD When i try i can be

OpenStudy (anonymous):

dang it i had just got the answer too >_< lold

OpenStudy (jadeishere):

xD

OpenStudy (anonymous):

only a few more questions till i get the grade! you guys are basically checking my work for me :D

OpenStudy (anonymous):

XD

OpenStudy (jadeishere):

checking? pish posh

OpenStudy (anonymous):

choices are: 1/4 -4 4

OpenStudy (anonymous):

i ended up with 1/4, but im pretty sure im wrong

OpenStudy (jadeishere):

it's 4

OpenStudy (anonymous):

i got 4

OpenStudy (anonymous):

see what i mean?

OpenStudy (jadeishere):

Lol >.<

OpenStudy (anonymous):

lol XD

OpenStudy (anonymous):

choices are: 19xsqrt3x 5xsqrt3x 13xsqrt3x

OpenStudy (jadeishere):

5x sqrt3x

OpenStudy (anonymous):

i got b

OpenStudy (anonymous):

thnx

OpenStudy (jadeishere):

yep >.<

OpenStudy (anonymous):

sqrt8(sqrt2 - x) = 11 solve for x

OpenStudy (anonymous):

brb

OpenStudy (anonymous):

kk jacob

OpenStudy (jadeishere):

\[\sqrt{8}\sqrt{2 - 3}\]

OpenStudy (jadeishere):

is that what it?

OpenStudy (anonymous):

no...

OpenStudy (jadeishere):

Okay then i don't understand

OpenStudy (jadeishere):

oh i just saw the = 11 /).<

OpenStudy (anonymous):

first option

OpenStudy (anonymous):

second option

OpenStudy (jadeishere):

\[\sqrt{8} \sqrt{2 - x} = 11\]is that it?

OpenStudy (anonymous):

third option

OpenStudy (anonymous):

solving for X

OpenStudy (jadeishere):

I don't understand what t looks like in the beginning so i have no clue what to start with

OpenStudy (anonymous):

\[\sqrt{8}\left( \sqrt{2-x} \right)=11\]

OpenStudy (anonymous):

thats the problem. have to solve for X

OpenStudy (jadeishere):

I honestly don't know O.O Sorry

OpenStudy (anonymous):

s'ok.

OpenStudy (anonymous):

\[(10x ^{6}+20x ^{4}-15x ^{2})\div 5x ^{2}\]

OpenStudy (anonymous):

i keep coming up with 3x^6

OpenStudy (anonymous):

but its not the correct answer

OpenStudy (anonymous):

it has empty spaces to fill in. ?x^4 + ?x^? - ?

OpenStudy (anonymous):

question marks are what is missing.

OpenStudy (jadeishere):

Can you take a screen shot please? That made no sense the way you worded it

OpenStudy (anonymous):